OpenStudy (anonymous):

Atmospheric Pressure at sea level is 1 atmosphere (atm). Under water pressure increases by 0.0303 atm for every 1 foot. Example water pressure in ocean is 1.0303 at a depth of 1 foot. let F(d) be water pressure at a water depth of d feet. What is the slope? I'm confused whether it is 0.0303 or 1? Thanks!

OpenStudy (anonymous):

F(d) = 1 + ).0303 d where units are atmospheres and feet y = mx + b = (slope) x + y-intercept form

OpenStudy (anonymous):

Yes, f(d) = 1+ 0.303d, where d is the number of feet.

OpenStudy (anonymous):

so 0.0303 is the slope?

OpenStudy (anonymous):

f(d) is y the slope is 0.3030, and one is the y-intercept. You can also write it as y = 0.0303x+1 y =mx+b

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

*0.0303

OpenStudy (anonymous):

Very sorry, typo, meant to write F(d) = 1 + 0.0303 d